OpenStudy (anonymous):

An object will take t sec to drop d feet, where t = \[\sqrt{d/16}\]. what distance will it drop in \[\sqrt{19}/2\] sec? I got 19ft.

OpenStudy (anonymous):

i think it might be 38 though

OpenStudy (lollylau):

76?

OpenStudy (paxpolaris):

\[t= \sqrt{d/16}={\sqrt{d}\over4}\]

OpenStudy (lollylau):

sqrt(19)/2=sqrt(d/16) 19/4=d/16 d=76?

OpenStudy (paxpolaris):

\[{\sqrt{19}\over2}={\sqrt{d}\over4}\]\[\sqrt d = 2 \sqrt{19}\] \[\huge \therefore d=76\]

OpenStudy (anonymous):

But the question isn't asking the value of d

OpenStudy (paxpolaris):

yes it is ... 76 ft

OpenStudy (lollylau):

d = distance distance = unknown unknown = answer therefore, d = answer

OpenStudy (anonymous):

Yes, answer is 76 feet

OpenStudy (anonymous):

either way i got it wrong thanks for all your help. i got 2 more i want to check but im too tired to type. gonna play a video game, hehe.